Before 3 years father age was 7 times of sons age after 3years father age will be 4 times of sons age calculate the present age sons and father
Answer by ankor@dixie-net.com(22740)

let f = father's age
let s = son's age
:
Write an equation for each statement, simplify
;
Before 3 years father age was 7 times of sons
f - 3 = 7(s-3)
f - 3 = 7s - 21
f = 7s - 21 + 3
f = 7s - 18
:
after 3years father age will be 4 times of sons
f + 3 = 4(s+3)
f + 3 = 4s + 12
f = 4s + 12 - 3
f = 4s + 9
:
f=f, so we can write this
7s - 18 = 4s + 9
7s - 4s = 9 + 18
3s = 27
s = 27/3
s = 9 yrs is the son
:
You can use f = 4s + 9, to find the father's age
:
Check your solutions in both statements.